OpenStudy (anonymous):

Solve x 2 ≤ 6 − x and sketch the solution on a number line 1. None of these 2. 3 ≤ x ≤ 2 3. −3 ≤ x ≤ 2 4. 0 ≤ x ≤ 1 5. 3 ≤ x ≤ −2 6. −3 ≤ x ≤ 0

OpenStudy (anonymous):

*x^2

OpenStudy (anonymous):

*x^2≤ 6 − x

OpenStudy (anonymous):

Hint: \[x^2+x-6=0\implies(x+3)(x-2)=0.\]

OpenStudy (anonymous):

what ? I dont get it

OpenStudy (anonymous):

The roots of this polynomial are at \(x=-3\) and \(x=2\). However, the inequality has it that \((x+3)(x-2)\leq0\). Therefore, picking a value inside the range \(-3\leq x\leq2\) will yield a negative number (prove this to convince yourself of it).